OpenStudy (imnotgoodatmath):

During a hike, Vijay saw a total of 19 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 52 wings. Use a model to determine how many birds he saw. 9 10 11 12

OpenStudy (imnotgoodatmath):

@ganeshie8

OpenStudy (ybarrap):

$$ \large 19=birds+butterflies $$ $$ 52=2*birds+4*butterflies $$ Can you solve this?

OpenStudy (chmvijay):

vijay!!!! you mean myself :) :)

OpenStudy (imnotgoodatmath):

:D

OpenStudy (abmon98):

Bf=Butterflies B-Birds B+Bf=19 2B+4Bf=52 Solve simultaneous equations, subtract both equations from each other. 2B+2Bf=38 2B+4Bf=52 -2Bf=38-52 Bf=14/2=7 19-7=12 Bird

OpenStudy (imnotgoodatmath):

so it would be 12?

OpenStudy (abmon98):

yes

OpenStudy (abmon98):

each butterfly has 4 wings and each bird has 2 wings, now we dont know the number of birds or butterflies Vijay saw but we know the total number of wings so now we have 2 unknowns. we represent them with variables or letters and solve it simultaneously with the total number of birds and butterflies seen by vijay. 2*(Number of birds)+4*(Number of Butterflies)=total number of wings. 2*(Number of birds)+4*(Number of Butterflies)=52